# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ IV Alleviating the effect of barren plateaux in variational quantum algorithms
On problems with a moderate size of optimization variables, VGON has shown remarkable ability to quickly converge to the (nearly) optimal output distribution and generate high quality solutions with high probability. In near-term hybrid quantum-classical algorithms such as the variational quantum eigensolver (VQE) [47], however, the number of classical parameters can quickly reach thousands or tens of thousands. The performance of such a hybrid algorithm can be hard to predict. On the classical part, the problems of vanishing gradients and having multiple minima are often present [17, 32, 39, 8, 46]. On the quantum part, the choice of ansätze greatly affects the expressivity of the quantum circuit, making the certification of global optimality difficult [27, 33, 51, 57, 7].
@@
where $\sigma_{x,y,z}^{i}$ denote the Pauli operators at site $i$. The ansatz for the parameterized quantum circuit is inspired by the matrix product state encoding [49]. It consists of sequential blocks of nearest-neighbor unitary gates, each of which is made of 11 layers of single qubit rotations and CNOT gates (see Appendix C.2 for more details).
By choosing $N=18$, the circuit contains $816$ blocks and $12,240$ variational parameters. The average ground state energy, computed using exact diagonalization (ED), is -1.7828. We use both VQE-SA and VGON to compute the same quantity, with each method repeated 10 times. The results are shown in Fig. 3, where the mean values and the 95% confidence intervals of these methods are visualized. The dark-blue and the green lines represent the average energy for VQE-SA and VGON, whose mean values at the last iteration are -1.7613 and -1.7802, respectively. Furthermore, to compare the performance of the two methods in a more fine-grained manner, we also calculate the fidelity between the states produced by the quantum circuit and the exact ground state. As the purple line depicted, VGON can achieve a 99% fidelity by around 880 iterations, while the VQE-SA method can only achieve 78.25% fidelity within the same number of iterations. Moreover, because VGON benefits from batch training, it leads to a more stable convergence. For another comparison, VQE with uniformly random initial parameters can barely provide meaningful results due to the presence of barren plateaux, where the mean value of the average energy is -0.1367 after 1,000 iterations across 10 repetitions, as illustrated by the light-blue line in Fig. 3.
Figure 3: Mean values and 95% confidence intervals of the energy densities and the fidelity to the exact ground state at different iterations. The light-blue line shows the average energy at different iterations for VQE. The dark-blue (red) and the green (purple) lines represent the average energies (fidelity between the produced state and the exact ground state) at different iterations for the VQE-SA method and VGON, respectively. The exact average ground state energy is depicted by the black dots.
@@ V Identifying degenerate ground state space of quantum models
Deterministic gradient-based optimization methods are predisposed to follow a single path, therefore hampering their ability to efficiently detect multiple optima. A unique advantage of generative models is the ability to produce diverse samples of objects, all of which may minimize objective function. In optimization, this leads to the possibility of finding multiple optimal solutions with a single stage of training. We now show that when appropriately trained, VGON exemplifies such an effective capability for generating multiple (nearly) optimal solutions simultaneously. This capability can be largely ascribed to its integration of randomness and the adoption of batch training. The former facilitates broader exploration within the variational manifold, while the latter, which involves processing subsets of data samples concurrently, supports the collective identification of multiple optimal solutions.
@@
where $\bm{\sigma}^{i}=(\sigma^{i}_{x},\sigma^{i}_{y},\sigma^{i}_{z})$ are Pauli operators at site $i$. We take $N=10$ for $H_{MG}$, and $N=11$ for $H_{232}$, making their ground state spaces 5- and 2-fold degenerate, respectively. An orthonormal basis for their respective degenerate ground state spaces is computed by the ED method, which outputs five vectors |$\ket{v_{1}}\ldots\ket{v_{5}}$ spanning the ground state space of $H_{MG}$, and two vectors |$\ket{u_{1}}$ and |$\ket{u_{2}}$ spanning that of $H_{232}$.
The overall objective of this problem is similar to the previous one: finding the ground states of $H_{MG}$ and $H_{232}$ with variational quantum circuits. We maintain the same circuit layout as in the previous problem, and use 36 and 60 blocks of unitary gates for each Hamiltonian respectively. Profiting from the use of mini-batches to estimate gradients, a common technique in training neural networks, VGON can effectively evaluate many different circuits simultaneously. Meanwhile, to enhance intra-batch diversity, a penalty term consisting of the mean cosine similarity among all pairs of sets of circuit parameters in the same batch is added to the objective function. This penalty term, together with the mean energy of the states in the batch, ensures a balance between maintaining the diversity of the generated outputs and minimizing the energy. Further details can be found in Appendix D.1.
@@
We generate 1,000 output states for each Hamiltonian using a VGON model trained with the above objective function. We find that the vast majority of these states have energy low enough to be treated as ground states. Figure 4 shows the overlap between the generated states and the basis of their ground state space. In Figure 4(a), the generated states for $H_{232}$ fall into two orthogonal classes, which form an orthonormal basis of the ground state space. For $H_{MG}$, Figure 4(b) shows that most of them are linearly independent and span the same space as |$\ket{v_{1}}\ldots\ket{v_{5}}$.
Figure 4: The overlap between 1,000 states generated by the trained VGON models and the orthonormal bases of the ground space computed by ED: |$\ket{u_{1}}$ and |$\ket{u_{2}}$ for $H_{232}$ shown in (a) and $|v_{1}\rangle,|v_{2}\rangle,\cdots,|v_{5}\rangle$ for $H_{MG}$ shown in (b). The shaded curves show the population densities of the generated states having different overlaps with one of the basis states.
VI Conclusions
@@
We propose a general approach called variational generative optimization network, or VGON for short, for tackling optimization challenges in a variety of quantum problems. This approach combines deep generative models in classical machine learning with sampling procedures and a problem-specific objective function, exhibiting excellent convergence efficiency and solution quality in quantum optimization problems of various sizes. Particularly, it alleviates the barren plateau problem, a pervasive issue in variational quantum algorithms, and surpasses the performance of the VQE-SA method, an approach designed specifically to avoid barren plateaux. Additionally, the capability of VGON to identify degenerate ground states of quantum many-body models underscores its efficacy in addressing problems with multiple optima. Due to its flexible design, we also envisage applying VGON to problems beyond the quantum world, particularly those in many areas of science and engineering where high quality multiple optimal solutions are expected.
Appendix A Reaching the quantum limit of a many-body contextuality witness
@@
where $\{\langle O_{x}^{1}\rangle:x\in\{a,b\}\}$, $\{\langle O_{x}^{1}O_{y}^{2}\rangle:x,y\in\{a,b\}\}$ and $\{\langle O_{x}^{1}O_{y}^{3}\rangle:x,y\in\{a,b\}\}$ are the expectations of single-site correlator, nearest-neighbor and next-to-nearest neighbor two-point correlators, respectively.
@@
For example, we can parameterize the local observables $O_{x}:x\in\{a,b\}$ as
@@
where $\Lambda_{x}$ is a diagonal matrix with entries being ±$\pm 1$, $\{S_{j}\}$ are the basis of the space of skew-symmetric matrices with the dimension of $m=(d^{2}-d)/2$, $d$ is the local dimension, and ${\bm{\theta}_{x}}\equiv(\theta_{x1},\theta_{x2},\dots,\theta_{xm})$ are real scalars. All the parameters combined are denoted as ${\bm{\theta}}\equiv(\bm{\theta}_{a},\bm{\theta}_{b})$. Then the Hamiltonian corresponding to contextuality witness (4) can be expressed as
@@
In fact, by a modified version of the Navascués-Pironio-Acín (NPA) hierarchy [65], a lower bound for the lowest GSED of $H(\bm{\theta})$ has been obtained to be -4.4142134689. However, whether there is any infinite TI quantum many-body Hamiltonian can achieve this bound is still unknown. Using stochastic gradient descent (SGD), Ref. [65] reports an infinite TI model that the corresponding GSED is $-4.4142131947$, which has a physical dimension $d=5$ and a bond dimension $D=5$.
@@
Appendix B Finding the optimal state for entanglement detection
@@
subject to ${\rm tr}(M_{N}\rho)=p_{2},\ p_{1}\mathbb{I}-M_{Y}\in\mathcal{S}^{*},$
$M_{Y}=M_{Y}(P),\ M_{N}=M_{N}(P),$
@@
B.1 The pure state case
@@
Here $U_{A}$ ($U_{B}$) can be parameterized by a set of $93^{2}=$ 9 linearly-independent skew-Hermitian matrices $\{T_{j}\}$ [25], i.e.,
@@
Here we set $p_{1}=({\rm tanh}(e_{1})+1)/2$, and ${p_{2}^{LO}}^{*}$ and $\ {p_{2}^{1-LOCC}}^{*}$ are computed by solving the two SDPs in Eq. (8) respectively.
To fully test the performance of SGD on this problem, we run this algorithm for 79,663 different initial states, which are sampled from the parameter space according to the distribution $\mathcal{N}(\bm{0,I})$. The results are listed in Fig. 5(a), where the green and the blue dots present the gaps for the initial states and the optimized states, respectively. It turns out that GD gets stuck in local minima easily, and only 1.52% of the initial states achieve a gap larger than 0.08. The largest gap observed is 0.0837. Lastly, we would like to stress that the above calculations take more than two months on a desktop-grade computer.
Figure 5: Distributions of gaps achieved by SGD and VGON. (a) Distributions of gaps obtained by SGD in two months. The green and blue dots present the gaps achieved by the 79,663 initial states of SGD and those by the corresponding optimized states, respectively. 1.52% of them achieve a gap larger than 0.08, which is represented by the dark-blue dotted line. The obtained maximal gap is roughly 0.0837. (b) Distributions of the gaps optimized by VGON in two hours. Remarkably, 98.59% of them exceed 0.08, and 52.657% of them even fall within the range $[0.0836,0.0837]$.
Subsequently, the same problem is addressed by VGON. The architecture of the VGON model consists of a 3-layer encoder network with sizes [512, 256, 128], a 3-layer decoder network with sizes [128, 256, 512], and a 2-dimensional latent space connecting the encoder and decoder components. The training is conducted with a batch size of 6, and an exponential decaying learning rate $lr_{i}$ at iteration $i$, where $lr_{i}=0.99\times lr_{i-1}$, and $lr_{1}=0.001$. Once trained, the VGON model generates 100,000 quantum states as the output, and the gaps achieved by these quantum states are listed in Fig. 5(b), where we can see that the performance of VGON overwhelmingly surpasses that of the GD method. Specifically, 98.59% of the quantum states that VGON generates exhibit a gap larger than 0.08, and 52.657% of them even fall within the range $[0.0836,0.0837]$, which is also the maximal gap found by SGD.
B.2 The mixed state case
@@
where $\Sigma$ is a $9\times 9$ diagonal matrix whose diagonal entries are nonnegative and sum to $1$, and $U$ is a unitary matrix that can be parameterized by a set of $19^{2}-$ 1 generalized Gell-Mann matrices $\{T_{j}\}$ [2], i.e.,
@@
where $\lambda_{j}$’s are $19^{2}-$ 1 real numbers, denoted as $\bm{\lambda}\in\mathbb{R}^{9^{2}-1}$.
@@
  • Parameter space: $\{e_{1}\in\mathbb{R},\bm{\lambda}\in\mathbb{R}^{9^{2}-1},\bm{\sigma}\in\mathbb{R}^{9}:\|\bm{\sigma}\|_{2}=1\}$
For such a task, the chosen VGON model comprises a 4-layer encoder network and a 3-layer decoder network with sizes [1024, 512, 256, 128] and [128, 256, 512] respectively. We maintain the same latent space dimension and the same learning rate as those used in the training for pure states. In addition, we train the VGON model with a batch size of 3. Our results are depicted in Fig. 6(a), which indicates that the VGON model is exceptionally adept at performing this task. Notably, 99.98% of the parameter sets generated by the VGON model manifest a gap of 0.07, and 99.32% of them even surpass a gap of 0.08. Particularly, as shown in Fig. 6(b), when starting from a variational submanifold of the space of all mixed states, VGON always identified an almost pure state maximizing the gap, where the minimum achieved purity is 0.9993, and 96.49% of the states have a purity greater than 0.9999. This shows the excellent capability of VGON in identifying qualified quantum states from the complex quantum state landscape without being stuck in local minima.
Figure 6: Numerical results for exhibiting the advantage of 1-LOCC over LO. (a) Distribution of the advantage brought by the VGON. (b) Distributions of the maximal eigenvalues for random training states and those generated by the VGON.

### v2

@@ IV Alleviating the effect of barren plateaux in variational quantum algorithms
On problems with a moderate size of optimization variables, VGON has shown its ability to quickly converge to the (nearly) optimal output distribution and generate high quality solutions with high probability. In near-term hybrid quantum-classical algorithms such as the variational quantum eigensolver (VQE) [58], however, the number of classical parameters can quickly reach thousands or tens of thousands. The performance of such a hybrid algorithm can be hard to predict. On the classical part, the problems of vanishing gradients and having multiple minima are often present [21, 38, 46, 10, 56]. On the quantum part, the choice of ansätze greatly affects the expressivity of the quantum circuit, making the certification of global optimality difficult [33, 40, 64, 70, 9].
@@
where $\sigma_{x,y,z}^{i}$ denote the Pauli operators at site $i$. The ansatz for the parameterized quantum circuit is inspired by the matrix product state encoding [61]. It consists of sequential blocks of nearest-neighbor unitary gates, each of which is made of 11 layers of single qubit rotations and CNOT gates (see Appendix The Heisenberg XXZ model for more details).
By choosing $N=18$, the circuit contains $816$ blocks and $12,240$ variational parameters. The average ground state energy, computed using exact diagonalization (ED), is -1.7828. We use both VQE-SA and VGON to compute the same quantity, with each method repeated 10 times. The results are shown in Table 2 and Figure 3, where the mean values and the 95% confidence intervals of these methods are visualized. The dark-blue and the green lines represent the average energy for VQE-SA and VGON, whose mean values at the last iteration are -1.7613 and -1.7802, respectively. Furthermore, to compare the performance of the two methods in a more fine-grained manner, we also calculate the fidelity between the states produced by the quantum circuit and the exact ground state. As the purple line depicted, VGON can achieve a 99% fidelity by around 880 iterations, while the VQE-SA method can only achieve 78.25% fidelity within the same number of iterations. We would like to remark that since the quantum computational resource consumed in VGON is similar to that in VQE-SA and classical computational resource is relatively cheap, the wall clock time cost per iteration for VGON can be comparable to that of VQE-SA. Moreover, the batch training of VGON can lead to a more stable convergence. For practical on-hardware quantum optimization, batch evaluation also enhances noise robustness by averaging over multiple circuit executions, thereby mitigating stochastic fluctuations, suppressing outliers, and accelerating convergence. For another comparison, VQE with uniformly random initial parameters can barely provide meaningful results due to the presence of barren plateaux, where the mean value of the average energy is -0.1367 after 1,000 iterations across 10 repetitions, as illustrated by the light-blue line in Figure 3.
Figure 3: Mean values and 95% confidence intervals of the energy densities and the fidelity to the exact ground state at different iterations. The light-blue line shows the average energy at different iterations for variational quantum eigensolver (VQE). The dark-blue (red) and the green (purple) lines represent the average energies (fidelity between the produced state and the exact ground state) at different iterations for the variational quantum eigensolver with small-angle initialization (VQE-SA) method and Variational Generative Optimization Network (VGON), respectively. The exact average ground state energy is depicted by the black dots. The inset zooms in on the convergence behavior of the average energies for VGON and VQE-SA, showcasing the faster convergence of VGON. Each method is repeated 10 times to calculate the mean values and 95% confidence intervals.
Table 2: Comparison between variational quantum eigensolver (VQE), VQE with small-angle initialization (VQE-SA), and Variational Generative Optimization Network (VGON).
- Optimal Mean
VQE VQE-SA VGON VQE VQE-SA VGON
Average Energy -0.1374 -1.7684 -1.7821 -0.01367 -1.7613 -1.7802
Fidelity - 90.00% 99.82% - 80.01% 99.17%
@@ V Identifying degenerate ground state space of quantum models
Deterministic gradient-based optimization methods are predisposed to follow a single path, therefore hampering their ability to efficiently detect multiple optima. A unique advantage of generative models is the ability to produce diverse samples of objects, all of which may minimize the objective function. In optimization, this leads to the possibility of finding multiple optimal solutions with a single stage of training. We now show that when appropriately trained, VGON exemplifies such an effective capability for generating multiple (nearly) optimal solutions simultaneously. This capability can be largely ascribed to its integration of randomness and the adoption of batch training. The former facilitates broader exploration within the variational manifold, while the latter, which involves processing subsets of data samples concurrently, supports the collective identification of multiple optimal solutions.
@@
[…]
The overall objective of this problem is similar to the previous one: finding the ground states of $H_{MG}$ and $H_{232}$ with variational quantum circuits. We maintain the same circuit layout as in the previous problem, and use 36 and 60 blocks of unitary gates for each Hamiltonian respectively. Profiting from the use of mini-batches to estimate gradients, a common technique in training neural networks, VGON can effectively evaluate many different circuits simultaneously. Meanwhile, to enhance intra-batch diversity, a penalty term consisting of the mean cosine similarity among all pairs of sets of circuit parameters in the same batch is added to the objective function. This penalty term, together with the mean energy of the states in the batch, ensures a balance between maintaining the diversity of the generated outputs and minimizing the energy. Further details can be found in Appendix The $Z_{1}Z_{2}$ model.
@@
[…]
VI Details
VI.1 Details on VGON
A VGON model contains two neural networks, i.e., the encoder $E_{\bm{\omega}}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{Z}$ and the decoder $D_{\bm{\phi}}\mathrel{\mathop{\ordinarycolon}}\mathcal{Z}\rightarrow\mathcal{X}$ connected by a latent space $\mathcal{Z}$, and they are parameterized by $\bm{\omega,\phi}$, respectively. These parameters are iteratively updated to produce a solution distribution $P_{\bm{\phi,\omega}}(\bm{x})$ such that the expectation of an objective function $h(\bm{x})$ is optimized:
$\mathbb{E}_{\bm{x}\sim P_{\bm{\phi,\omega}}(\bm{x})}\left[h(\bm{x})\right]$
$=$ $\int\int h\left(D_{\bm{\phi}}(\bm{z})\right)P_{\bm{\omega}}(\bm{z}|\bm{x_{0}})P(\bm{x_{0}})d\bm{x_{0}}d\bm{z}$
$=$ $\int h\left(D_{\bm{\phi}}(\bm{z})\right)P_{\bm{\omega}}(\bm{z})d\bm{z}$
$=$ $\mathbb{E}_{z\sim P_{\bm{\omega}}(\bm{z})}\left[h\left(D_{\bm{\phi}}(\bm{z})\right)\right].$
More specifically, the input data $\bm{x_{0}}$ is sampled from a given distribution $P(\bm{x_{0}})$, which is then mapped to the latent distribution $P_{\bm{\omega}}(\bm{z})$ by the encoding estimator $E_{\bm{\omega}}$, i.e., $P_{\bm{\omega}}(\bm{z})=\int P_{\bm{\omega}}(\bm{z}|\bm{x_{0}})P(\bm{x_{0}})d\bm{x_{0}}$. Next the decoding estimator $D_{\bm{\phi}}(\bm{z})$ further maps the latent distribution $P_{\bm{\omega}}(\bm{z})$ to a distribution $P_{\bm{\phi,\omega}}(\bm{x})$ of the target data $\bm{x}$, which is right the input of the objective function $h(\bm{x})$.
Additionally, for the convenience of training, we constraint the distribution of the latent space to be a normal distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$, and try to minimize the distance between it and the standard normal distribution $\mathcal{N}(0,I)$, measured by the KL divergence. Specifically, the cost function for VGON is formulated as
$C(\bm{\phi,\omega})$ (4)
$=$ $\mathbb{E}_{\bm{x}\sim P_{\bm{\phi,\omega}}(\bm{x})}\left[h(\bm{x})\right]+\beta\cdot\mathcal{D}\left[\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})||\mathcal{N}(0,I)\right],$
where the hyperparameter $\beta$ represents the trade-off between the expectation of the objective function and the above KL divergence.
In our implementations of VGON, all the training procedures are conducted based on the PyTorch framework [57]. To address different tasks, diverse objective functions $h(\bm{x})$ are employed, and each requires a specific interfacing with PyTorch. The configurations of these VGON models will be detailed in the subsequent sections, which can provide us a comprehensive understanding of how VGON is tailored for different optimization challenges.
VI.2 Datails on finding the optimal states in entanglement detection
To find the quantum state that can exhibit the largest advantage of 1-LOCC protocols over LO protocols in entanglement detection, the problem can be formulated as maximizing the difference between the solutions to the two SDPs introduced in Eq. (1), with the following objective function and parameter space:
  • Objective function: $h(\rho^{\prime},p_{1})={p_{2}^{LO}}^{*}(\rho^{\prime},p_{1})-{p_{2}^{1-LOCC}}^{*}(\rho^{\prime},p_{1})$
  • Parameter space: $\{e_{1}\in\mathbb{R},\rho^{\prime}\in\{\rho_{exp}\}\text{ or }\{\rho\}\}$
Here, $p_{1}$ is parameterized as $p_{1}=({\rm tanh}(e_{1})+1)/2$, and $\{\rho_{exp}\}$ and $\{\rho\}$ represent the set of states for the pure case and mixed case, respectively.
Efficient variational optimization for these SDPs and their integration into the PyTorch framework for machine learning requires the use of CVXPY and cvxpylayers [13, 2, 1]. The first translates a convex optimization problem into a form that solvers can understand, while the latter allows automatic differentiation of convex optimization problems by computing their gradients and backpropagate them through the neural network.
As we mentioned in the main text, the state space we consider follows closely the linear optical setup which generates arbitrary bipartite qutrit states $\rho_{\rm exp}$. Photons generated by the laser source are expressed as $\sum_{i}c_{i}|i\rangle$, where $c_{i}$ are complex numbers satisfying $\sum_{i}|c_{i}|^{2}=1$. Afterwards, these photons go through spontaneous parametric down-conversion (SDPC), which converts their state to $|\psi\rangle=\sum_{i}c_{i}|ii\rangle$. In the case of qutrits, the state can be parameterized as [28, 27]
where $\phi,m,n\in[0,2\pi)$, and $\theta\in[0,\pi]$ are variational parameters. Subsequently, two local unitaries denoted by $U_{A},U_{B}$ can be applied on the two subsystems, resulting in the quantum state
Here, $U_{A}$ ($U_{B}$) can be parameterized by a set of $3^{2}=9$ linearly-independent skew-Hermitian matrices $\{T_{j}\}$ [29], i.e.,
$U_{A}\ (U_{B})=\exp\left(\sum_{j=1}^{9}\lambda_{j}T_{j}\right),$
where $\lambda_{j}$’s are $9$ real numbers, denoted as $\bm{\lambda}_{A}$ ($\bm{\lambda}_{B}$). Therefore, the parameterized space for pure states $\rho_{\rm exp}$ is represented as
$\{m,n,\phi,\theta\in\mathbb{R},\ \bm{\lambda}_{A},\bm{\lambda}_{B}\in\mathbb{R}^{9}\}.$
On the other hand, mixed qutrit states $\rho\in\mathcal{H}^{3}\otimes\mathcal{H}^{3}$ can be parameterized by
$\rho=U\Sigma U^{\dagger},$
where $\Sigma$ is a $9\times 9$ diagonal matrix whose diagonal entries are nonnegative and sum to $1$, and $U$ is a unitary matrix that can be parameterized by a set of $9^{2}-1$ generalized Gell-Mann matrices $\{T_{j}\}$ [5], i.e.,
$U=\exp\left(i\sum_{j=1}^{9^{2}-1}\lambda_{j}T_{j}\right),$
where $\lambda_{j}$’s are $9^{2}-1$ real numbers, denoted as $\bm{\lambda}\in\mathbb{R}^{9^{2}-1}$. Furthermore, the normalized diagonal matrix $\Sigma$, denoted as ${\rm diag}(\sigma_{1}^{2},\cdots,\sigma_{9}^{2})$, can be obtained by ensuring that the Euclidean norm of the vector $\bm{\sigma}$ is equal to 1, i.e., $|\bm{\sigma}|_{2}=1$, where $\bm{\sigma}=(\sigma_{1},\cdots,\sigma_{9})$. Consequently, the parameterized space for mixed states $\rho$ case is written as
$\{\bm{\lambda}\in\mathbb{R}^{9^{2}-1},\bm{\sigma}\in\mathbb{R}^{9}\mathrel{\mathop{\ordinarycolon}}|\bm{\sigma}|_{2}=1\}.$
VI.3 Details on alleviating barren plateaus in variational quantum algorithms
A typical VQE algorithm can approximate the ground state of a given Hamiltonian $H$ using a variational wave function generated by a parameterized quantum circuit (PQC) $U(\bm{\theta})$, represented as $|\psi(\bm{\theta})\rangle=U(\bm{\theta})|00\cdots 0\rangle$. This sets up the minimization problem:
  • Parameter space: $\{\bm{\theta}\in\mathbb{R}^{M}\}$
The dimension of the parameter space $M$ is determined by the structure of the PQC.
The simulation of PQCs and the computation of energy are implemented by PennyLane [4], a software library for quantum machine learning. Its support of the CUDA-based CuQuantum SDK from NVIDIA enables VGON to handle over 10000 variational parameters on a consumer grade graphics card. PennyLane also provides seamless integration with PyTorch and its machine learning toolkit. Efficient GPU-accelerated simulation of PQCs is achieved by using the adjoint differentiation method [31] to compute the gradients, after which the parameters are updated by the Adam optimizer.
One of the key differences between VQE, VQE-SA and VGON is the initialization of parameters. For the VQE and VQE-SA, the initial parameters $\bm{\theta}$ are uniformly sampled from the range $0,2\pi)$ and $[0,0.01)$, respectively. In VGON, on the other hand, the decoder initialized using PyTorch’s default settings generates the circuit’s initial parameters $\bm{\theta}$. For more details on these methodologies and the comparisons between their performance, please refer to the third subsection in the Results and Appendix Finding the optimal state for entanglement detection.
VI.4 Details on identifying degenerate ground state space of quantum models
where $\mathcal{C}_{S_{b}}^{2}$ represents the set of all 2-combinations pairs derived from the elements in $\{1,2,\cdots,S_{b}\}$, and $\|\cdot\|$ denotes the Euclidean norm. Eventually, the optimization objective is set as minimizing the linear combination of $\bar{E}(\bm{\Theta})$ and $\bar{S}_{\mathcal{C}_{S_{b}}^{2}}(\bm{\Theta})$ according to a trade-off coefficient $\gamma$, i.e.,
  • Parameter space: $\{\bm{\Theta}\in\mathbb{R}^{S_{b}M}\}$
We estimate the quality of the generated state by computing the overlap between them and a basis of the ground space. Such computations can be resource-intensive, and therefore we only demonstrate the performance of VGON for 10-qubit systems.
VII Conclusions
@@
We propose a general approach called variational generative optimization network, or VGON for short, for tackling variational optimization challenges in a variety of quantum problems. This approach combines deep generative models in classical machine learning with sampling procedures and a problem-specific objective function, exhibiting excellent convergence efficiency and solution quality in quantum optimization problems of various sizes. Particularly, it may alleviate the barren plateau problem, a pervasive issue in variational quantum algorithms, and surpasses the performance of the VQE-SA method, an approach designed specifically to avoid barren plateaux. Additionally, the capability of VGON to identify degenerate ground states of quantum many-body models underscores its efficacy in addressing problems with multiple optima. Beyond the quantum world, generative models are emerging as powerful tools in the field of optimization problems. For instance, diffusion models are now being utilized for combinatorial optimizations [67]. Due to the flexible designs, we also envisage VGON and such algorithms to complement each other in addressing a broader spectrum of optimization challenges.
Acknowledgements This work is supported by the Sichuan Provincial Key R&D Program (2024YFHZ0371), the National Natural Science Foundation of China (62250073, 62272259, 62332009) and the National Key R&D Program of China (2021YFE0113100). The authors would like to thank Abolfazl Bayat, Dongling Deng, Chu Guo, Zhengfeng Ji, Damian Markham, Miguel Navascués and Ying Tang for helpful comments.
Code availability The complete code of this study is openly accessible via the GitHub repository https://github.com/zhangjianjianzz/VGON.
Appendix
Reaching the quantum limit of a many-body contextuality witness
@@
where $\{\langle O_{x}^{1}\rangle\mathrel{\mathop{\ordinarycolon}}x\in\{a,b\}\}$, $\{\langle O_{x}^{1}O_{y}^{2}\rangle\mathrel{\mathop{\ordinarycolon}}x,y\in\{a,b\}\}$ and $\{\langle O_{x}^{1}O_{y}^{3}\rangle\mathrel{\mathop{\ordinarycolon}}x,y\in\{a,b\}\}$ are the expectations of single-site correlator, nearest-neighbor and next-to-nearest neighbor two-point correlators, respectively.
@@
For example, we can parameterize the local observables $O_{x}\mathrel{\mathop{\ordinarycolon}}x\in\{a,b\}$ as
@@
where $\Lambda_{x}$ is a diagonal matrix with entries being $\pm 1$, $\{S_{j}\}$ are the basis of the space of skew-symmetric matrices with the dimension of $m=(d^{2}-d)/2$, $d$ is the local dimension, and ${\bm{\theta}_{x}}\equiv(\theta_{x1},\theta_{x2},\dots,\theta_{xm})$ are real scalars. All the parameters combined are denoted as ${\bm{\theta}}\equiv(\bm{\theta}_{a},\bm{\theta}_{b})$. Then the Hamiltonian corresponding to contextuality witness (5) can be expressed as
@@
In fact, by a modified version of the Navascués-Pironio-Acín (NPA) hierarchy [79], a lower bound for the lowest GSED of $H(\bm{\theta})$ has been obtained to be -4.4142134689. However, whether there is any infinite TI quantum many-body Hamiltonian can achieve this bound is still unknown. Using stochastic gradient descent (SGD), Ref. [79] reports an infinite TI model that the corresponding GSED is 4.4142131947-4.4142131947, which has a physical dimension $d=5$ and a bond dimension $D=5$.
@@
Finding the optimal state for entanglement detection
@@
subject to ${\rm tr}(M_{N}\rho)=p_{2},$
$p_{1}\mathbb{I}-M_{Y}\in\mathcal{S}^{*},$
@@
The pure state case
@@
Here $U_{A}$ ($U_{B}$) can be parameterized by a set of $3^{2}=9$ linearly-independent skew-Hermitian matrices $\{T_{j}\}$ [29], i.e.,
@@
Here we set $p_{1}=({\rm tanh}(e_{1})+1)/2$, and $p_{2}^{LO}$$^{*}$ and $p_{2}^{1-LOCC}$$^{*}$ are computed by solving the two SDPs in Supplementary Eq. (9) respectively.
To fully test the performance of SGD on this problem, we run this algorithm for 79,663 different initial states, which are sampled from the parameter space according to the distribution $\mathcal{N}(\bm{0,I})$. The results are listed in Supplementary Figure 1(a), where the green and the blue dots present the gaps for the initial states and the optimized states, respectively. It turns out that GD gets stuck in local minima easily, and only 1.52% of the initial states achieve a gap larger than 0.08. The largest gap observed is 0.0837. Lastly, we would like to stress that the above calculations take more than two months on a desktop-grade computer.
Supplementary Fig. 1: Distributions of gaps achieved by SGD and VGON. (a) Distributions of gaps obtained by SGD in two months. The green and blue dots present the gaps achieved by the 79,663 initial states of SGD and those by the corresponding optimized states, respectively. 1.52% of them achieve a gap larger than 0.08, which is represented by the dark-blue dotted line. The obtained maximal gap is roughly 0.0837. (b) Distributions of the gaps optimized by VGON in two hours. Remarkably, 98.59% of them exceed 0.08, and 52.657% of them even fall within the range $[0.0836,0.0837]$.
Subsequently, the same problem is addressed by VGON. The architecture of the VGON model consists of a 3-layer encoder network with sizes [512, 256, 128], a 3-layer decoder network with sizes [128, 256, 512], and a 2-dimensional latent space connecting the encoder and decoder components. The training is conducted with a batch size of 6, and an exponential decaying learning rate $lr_{i}$ at iteration $i$, where $lr_{i}=0.99\times lr_{i-1}$, and $lr_{1}=0.001$. Once trained, the VGON model generates 100,000 quantum states as the output, and the gaps achieved by these quantum states are listed in Supplementary Figure 1(b), where we can see that the performance of VGON overwhelmingly surpasses that of the GD method. Specifically, 98.59% of the quantum states that VGON generates exhibit a gap larger than 0.08, and 52.657% of them even fall within the range $[0.0836,0.0837]$, which is also the maximal gap found by SGD.
The mixed state case
@@
where $\Sigma$ is a $9\times 9$ diagonal matrix whose diagonal entries are nonnegative and sum to $1$, and $U$ is a unitary matrix that can be parameterized by a set of $9^{2}-1$ generalized Gell-Mann matrices $\{T_{j}\}$ [5], i.e.,
@@
where $\lambda_{j}$’s are $9^{2}-1$ real numbers, denoted as $\bm{\lambda}\in\mathbb{R}^{9^{2}-1}$.
@@
  • Parameter space: $\{e_{1}\in\mathbb{R},\bm{\lambda}\in\mathbb{R}^{9^{2}-1},\bm{\sigma}\in\mathbb{R}^{9}\mathrel{\mathop{\ordinarycolon}}\|\bm{\sigma}\|_{2}=1\}$
For such a task, the chosen VGON model comprises a 4-layer encoder network and a 3-layer decoder network with sizes [1024, 512, 256, 128] and [128, 256, 512] respectively. We maintain the same latent space dimension and the same learning rate as those used in the training for pure states. In addition, we train the VGON model with a batch size of 3. Our results are depicted in Supplementary Figure 2(a), which indicates that the VGON model is exceptionally adept at performing this task. Notably, 99.98% of the parameter sets generated by the VGON model manifest a gap of 0.07, and 99.32% of them even surpass a gap of 0.08. Particularly, as shown in Supplementary Figure 2(b), when starting from a variational submanifold of the space of all mixed states, VGON always identified an almost pure state maximizing the gap, where the minimum achieved purity is 0.9993, and 96.49% of the states have a purity greater than 0.9999. This shows the excellent capability of VGON in identifying qualified quantum states from the complex quantum state landscape without being stuck in local minima.
Supplementary Fig. 2: Numerical results for exhibiting the advantage of 1-LOCC over LO. (a) Distribution of the advantage brought by the VGON. (b) Distributions of the maximal eigenvalues for random training states and those generated by the VGON.
Comparison between VGON and various global optimization algorithms